OpenStudy (anonymous):

\[2x ^2\times 3x ^{7}/15x ^{6} (fraction problem)\]

OpenStudy (anonymous):

\[x^m\times x^n=x^{m+n}\] \[x^m\div x^n=x^{m-n}\]

OpenStudy (anonymous):

\[I see thanks Tomas. A. Also I have another problem I'm little confus\]e (2x ^{-3}yz ^{-6)(2x ^{?}}\]

OpenStudy (anonymous):

can you rewrite it? :D

OpenStudy (anonymous):

I see thanks Tomas. A. Also I have another problem I'm little confuse (2x ^{-3}yz ^{-6) (2x \[I see thanks Tomas. A. Also I have another problem I'm little confuse 2x ^{-3} yz ^{-6}) (2x)^{-5} \]

OpenStudy (anonymous):

i can't understand anyway, lol

OpenStudy (anonymous):

\[(2x ^{-3} yz ^{-6}) (2x)^{-5}\]

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

\[(ab)^n=a^nb^n\]

OpenStudy (anonymous):

thanks again Tomas. A:)